1. /16.66 points ASWSBE13 13.E.002.MI My Notes Ask Your Teacher You may need to use the appropriate technology to answer this question. In a completely randomized design, eight experimental units were used for each of the five levels of the factor. Complete the following ANOVA table. (Round your values for MSE and F to two decimal places, and your p-value to four decimal places.) Sum Source of Variation of Squares Degrees of Freedom Mean Square p-value 300 Error Total 480 Need Help? İtadeliL Master It Talk to a Tutor
Answer #1

N = 8*5 = 40

n = 8

Degress of Freedom:

dfBetween = a – 1 = 5-1 =4

dfWithin = N-a = 40-5 = 35

dfTotal = N-1 = 40-1 = 39

MS = Sum of Squares / degrees of Freedom

F = MSTreatments / MSError

Completed ANOVA Table

degrees of Sum of Squares FreedomMS p-value Treatments Error Total 300 180 480 4 35 39 75 14.583 <0.00001 5.14